Circular No. 8546

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 9P/TEMPEL D. Schleicher, Lowell Observatory; and K. Barnes, Franklin and Mar-

shall College, report that they obtained narrowband photometry of comet 9P on June 9 using the Hall 1.1-m telescope at Lowell Observatory, with the following averaged results: log Q(OH) = 27.79; equivalent log Q(water; vectorial) = 27.83; log Afρ = 2.1 (cf. IAUC 7342). No trends with aperture size are evident for the gas species, but the radial fall-off of the dust is significantly steeper than the canonical 1/ρ. When combined with monthly photometry previously obtained by Schleicher beginning in March, it is clear that water production is consistently about a factor of 2.4 lower than during the equivalent time interval in 1983, and CN is lower by about 2.0 times. Dust production, based on the proxy Afρ, is also lower, but only by about 1.3 times. Measurements from 1994 show intermediate values between the 1983 and 2005 results, most likely implying an on-going significant secular decrease in vaporization due to the exhaustion or covering over of a major source region on the nucleus.

SUPERNOVAE 2005bo AND 2005bt M. A. Pérez-Torres, M. Jelínek, J. Gorosabel, A. de Ugarte Postigo,

A. J. Castro-Tirado, A. Sota, and A. Alberdi, Instituto de Astrofísica de Andalucía (IAA), Consejo Superior de Investigaciones Científicas, Granada, report R-band magnitudes (± 0.1) for two supernovae obtained with the IAA 0.6-m and 1.5-m telescopes at Sierra Nevada Observatory: SN 2005bo in NGC 4708 (cf. IAUC 8514), Apr. 29.00 UT, 15.2; May 22.88, 16.0. SN 2005bt in UGC 8205 (cf. IAUC 8518), Apr. 29.10, 17.0.

VARIABLE OBJECT IN BOOTES R-band magnitudes for this object (cf. IAUC 8518) by Pérez-Torres et

al.: Apr. 29.07 UT, 16.5; May 18.94, 18.0.

COMET C/2000 S6 (SOHO) Following is the initial available position for another Kreutz sungrazing

comet found on SOHO website images (continuation to IAUC 8530). K. Battams adds that this object appeared stellar and was barely above the background noise.

Comet 2000 UT α_{2000} δ_{2000} Inst. F MPEC
C/2000 S6 Sept.28.488 12^{h}03.^{m}2 − 3^{o}03′ C2 HS 2005-K43
2005 June 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green